Minimize Swarm Operational News

Extension of Swarm 2 Hz TII Cross-Track data flow dataset

08 December 2017

We inform users that the 2 Hz TII Cross-Track data flow dataset has been extended. The data now covers 01 December 2015 to 29 April 2017 for the A and B satellites.

New Swarm acceleration and thermosphere density data released

24 November 2017

We are pleased to announce the new release of non-gravitational acceleration and thermosphere neutral density data, which were derived from the measurements of the GPS receiver. The newly released data cover the period covering the start of the mission until 19 July 2014. Hence, the available time series is now complete until 31 August 2017 for all three Swarm satellites.

Web Content Image

ESA Member States approve Swarm extension

20 November 2017

Timely placed a few days in advance of Swarm's fourth anniversary in space, ESA Member States participating to the Earth Observation Envelope Programme have given a nice birthday present to our mission.

Swarm MAGx_LR_1B dataset (version 0502) has been extended

06 November 2017

With reference to the news published on 07 September 2017, we inform the Swarm users that the Swarm MAGx_LR_1B version 0502 has been extended for all Swarm spacecraft.

Swarm Charlie MAGx_LR_1B vs.0502 of May 2017 regenerated

02 November 2017

We would like to inform Swarm users that due to the occurrence of an issue during the processing of the Swarm L1B MAGx_LR_1B data with Product Baseline and File Counter 0502, the Swarm Charlie VFM data for the time window from 01 to 31 May 2017 is erroneously set to zero.

Swarm Alpha MAGx_LR_1B of 29 April – 01 May 2017 now available

10 October 2017

With reference to the news published on 04 May 2017 we are pleased to inform the Swarm users that  the Swarm MAGA_LR_1B data for the days from 29 April to 01 May 2017 have been generated and now can be found in the ESA ftp server.

Publication of 16 Hz TII cross-track data flow dataset

29 September 2017

We inform the Swarm users that the 16 Hz TII Cross-Track data flow dataset has been made available in the ESA ftp server covering the time interval from 01/01/2016 to 29/04/2017 for S/C A and B.

US job opening on Swarm/e-POP

19 September 2017

The Department of Physics and Astronomy at the University of Iowa seeks to appoint a post-doctoral researcher to join a dynamic and expanding group of space physics researchers. Amongst other sources, the post-holder will be using Swarm mission data to aid in studies of the Earth's magnetic field.

New release of improved Swarm acceleration and thermosphere density data

18 September 2017

We are pleased to announce a new release of non-gravitational acceleration and thermosphere neutral density data, which was derived from the measurements of the GPS receiver.

Swarm Bravo: L1B and L2 Cat-2 data gap

14 September 2017

We inform the Swarm users that due to an anomaly that occurred on the VFM (Vector Field Magnetometer) on board Swarm Bravo, i.e., VFM generation time of science packet frozen from 11/09/2017 03:54 UTC to 13/09/2017 12:24 UTC, the Swarm Level 0 data for this time window was not generated.

Swarm ASM-VFM residual dataset extended

13 September 2017

With reference to the news published on 08 July 2016 we inform the Swarm users that the ASM-VFM Residual dataset produced for the investigations of the scalar residuals between the readings of the scalar magnetometer (ASM) and the vector magnetometer (VFM) has been extended until 31 August 2017 for all Swarm spacecraft.

Swarm MAGA_LR_1B data from 14 to 17 Dec 2013 now available

13 September 2017

We are pleased to inform the Swarm users that the 4-days gap present in the Swarm Alpha MAGx_LR_1B data almost at the beginning of the mission, i.e., from the 14 to 17 December 2013, has been recovered.

Improved version of the Swarm MAGx_LR_1B data now available

07 September 2017

We are pleased to inform the Swarm users that an improved version of the MAGx_LR_1B products (Product Baseline and File Counter 0502) is now available in the "Advanced" folder of the ESA FTP server. This dataset covers the time window from the beginning of the mission to end August 2017 for all Swarm spacecraft.

Orbit Maintenance Manoeuvre on Swarm Alpha on 13 July 2017

19 July 2017

We would like to inform the Swarm users that an Orbit Maintenance Manoeuvre was successfully performed on Swarm Alpha on 13 July 2017 starting at 14:17:55 UTC.

Changes introduced in the Swarm Dissemination FTP layout

12 July 2017

In relation to the previous communication we are pleased to inform you that the announced changes to the FTP layout are now effective as planned.

Absolute Vertical TEC now available in the Swarm Level 2 Cat-2 TEC products

10 July 2017

We inform the Swarm users that, starting from 05 July 2017, the Absolute Vertical TEC (VTEC) values and corresponding elevation angles will be available in the Swarm Level 2 Cat-2 TEC products.

Data gap present in Swarm-Charlie L1B production on 04 July 2017

10 July 2017

We inform the Swarm Users that because of an issue during the retrieval of the telemetry data from Swarm Charlie, a data gap is present in the Level 1B production on 04 July 2017 from 06:02 to 13:43 UTC.

Reminder of changes to be introduced in the Swarm Dissemination FTP layout

07 July 2017

Following the previous communication, we are pleased to inform you that the announced changes to the FTP layout will be effective on 12 July 2017. Please be aware that some service disruptions might happen during that day.

Changes to Swarm Dissemination FTP structure

27 June 2017

We are pleased to inform that, following a Swarm community recommendation, several changes will be introduced in the Swarm Dissemination (swarm-diss.eo.esa.int) FTP structure in early July 2017.

Swarm Alpha and Bravo data from June regenerated

19 June 2017

We inform the Swarm users that because of an issue during the processing of the Swarm L1B Orbit and Attitude product of 03/06/2017 on Swarm Alpha and on 08/06/2017 on Swarm Bravo - i.e., available Level 0 data was not included in in the L1B processing - the L1B data production has been affected.

Showing 1 - 20 of 139 results.
of 7